A gas made up of homonuclear diatomic molecules escapes through a pinhole 0.268 times as fast as H2 gas. Write the chemical formula of the gas


1
Expert's answer
2021-05-03T07:24:49-0400

"0.268=\sqrt{\frac {2} {M}}"


"0.5177=\frac {2} {M}"


"M=3.864=4"


Chemical formula of the gas;He